### v1

# Viscoelastic Worthington jets & droplets produced by bursting bubbles

Abstract
Bubble bursting and subsequent collapse of the open cavity at free surfaces of contaminated liquids can generate aerosol droplets, facilitating pathogen transport. After film rupture, capillary waves focus at the cavity base, potentially generating fast Worthington jets that are responsible for ejecting the droplets away from the source. While extensively studied for Newtonian fluids, the influence of non-Newtonian rheology on this process remains poorly understood. Here, we employ direct numerical simulations to investigate the bubble cavity collapse in viscoelastic media, such as polymeric liquids, examining how their elastic modulus $G$ and their relaxation time $\lambda$ affect jet and droplet formation. We show that the viscoelastic liquids yield Newtonian-like behavior as either parameter $G$ or $\lambda$ approaches zero, while increasing them suppresses jet formation due to elastic resistance to elongational flows. Intriguingly, for some cases with intermediate values of $G$ and $\lambda$, smaller droplets are produced compared to Newtonian fluids, potentially enhancing aerosol dispersal. By mapping the phase space spanned by the elastocapillary number (dimensionless $G$) and the Deborah number (dimensionless $\lambda$), we reveal three distinct flow regimes: (i) jets forming droplets, (ii) jets without droplet formation, and (iii) absence of jet formation. Our results elucidate the mechanisms underlying aerosol suppression versus fine spray formation in polymeric liquids, with implications for pathogen transmission and industrial processes involving viscoelastic fluids.
1 Introduction
Bubbles in liquids (Lohse, 2018) – from oceans (Deike, 2022) and volcanoes (Gonnermann & Manga, 2007) to cosmetic gels (Lin, 1970; Daneshi & Frigaard, 2024) and champagne (Liger-Belair, 2012; Mathijssen et al., 2023) – rise due to buoyancy and reach the liquid-gas interface, where they sit as the intervening liquid film drains (figure 1a-i, Lhuissier & Villermaux, 2012; Bartlett et al., 2023). Upon film rupture, numerous tiny droplets, known as film droplets, scatter over the free surface (Lhuissier & Villermaux, 2012; Villermaux et al., 2022), leaving a high-energy bubble cavity (figure 1a-ii, Woodcock et al., 1953; Knelman et al., 1954; Mason, 1954). The subsequent collapse of this cavity is driven by surface tension. This process involves rim retraction (Taylor, 1959; Culick, 1960; Sanjay et al., 2022) that generates capillary waves (Eggers et al., 2025). These waves propagate along the cavity, converging at its base to create an inertial flow focusing (Gordillo & Rodríguez-Rodríguez, 2019; Gordillo & Blanco-Rodríguez, 2023) that forms a Worthington jet (Worthington, 1877, 1908; Stuhlman Jr, 1932; Lohse et al., 2004; Sanjay, 2022) that features large strain rates (Sen et al., 2024). The jet may fragment into droplets through end-pinching and the Rayleigh–Plateau instability (Lord Rayleigh, 1878; Plateau, 1873; Keller et al., 1995; Stone & Leal, 1989; Ghabache & Séon, 2016; Walls et al., 2015). These jet droplets, typically larger and faster than the initial film droplets, play a crucial role in transporting dissolved substances to the atmosphere (Berny et al., 2020; Villermaux et al., 2022; Dubitsky et al., 2023a). The dynamics of bubble bursting have far-reaching implications across various domains. These include the transfer of pathogens from contaminated water to air (Bourouiba, 2021), the transport of dissolved salt from seawater to the atmosphere, where salt particles act as cloud condensation nuclei (Dubitsky et al., 2023b; De Leeuw et al., 2011), and the dynamics in bioreactors containing animal cells (Boulton-Stone & Blake, 1993). The unique capacity of ejected droplets to transport diverse species underscores the importance of comprehending the complete dynamics that dictate their formation. Ever since the first documented study of Stuhlman Jr (1932), advanced experiments and simulations have extensively characterized the rich dynamics of bursting bubbles. Key metrics include ejected drop heights (Stuhlman Jr, 1932), sizes (Kientzler et al., 1954; Deike et al., 2018; Berny et al., 2020, 2021; Blanco-Rodríguez & Gordillo, 2020; Villermaux et al., 2022), and velocities (Deike et al., 2018; Gordillo & Rodríguez-Rodríguez, 2019; Sanjay et al., 2021; Gordillo & Blanco-Rodríguez, 2023).
MacIntyre (1972) revealed internal liquid flow using dye and attempted to understand the drop composition, which was finally explained by direct numerical simulations (DNS) of Dubitsky et al. (2023a). Furthermore, Dasouqi et al. (2022) demonstrated atmospheric flow patterns using smoke-filled bubbles, which were detailed numerically by Singh & Das (2021). Although shadowgraphy techniques limit most experimental studies, x-ray imaging has captured traveling capillary wave dynamics, providing crucial validation for DNS results (Lee et al., 2011). These advancements have significantly enhanced our understanding of bubble bursting at the Newtonian liquid-gas interface across various scales and applications. Indeed, for a bubble of radius $R_{0}$ surrounded by a liquid with viscosity, density, and surface tension $\eta_{s}$, $\rho_{s}$, and $\gamma$, the interplay of capillarity, viscosity, and gravity governs the bubble cavity collapse. Correspondingly, the key control parameters of this process are the solvent Ohnesorge number
$Oh_{s}=\frac{\eta_{s}}{\sqrt{\rho_{s}\gamma R_{0}}},$ (1)
and the Bond number
$Bo=\frac{\rho_{s}gR_{0}^{2}}{\gamma}.$ (2)
Here, $g$ is the acceleration due to gravity. The solvent Ohnesorge number $Oh_{s}$ exemplifies the dimensionless viscosity of the surrounding medium, significantly influencing the capillary wave dynamics, determining their damping and overall viscous dissipation, while the Bond number $Bo$ affects the initial cavity shape and the hydrostatic pressure differences (Walls et al., 2015; Bergmann et al., 2006, 2009; Lohse, 2018). In this study, we will focus our attention on the limiting case of very small bubbles with $Bo=0.001$, for which the bubbles can be approximated as spheres (figures 1a, Toba, 1959; Princen, 1963; Lhuissier & Villermaux, 2012). For the Newtonian cases, appendix A summarizes the key results, including the effect of $Oh_{s}$ on bubble-busting dynamics. For the influence of gravity on the shape and consequently the overall dynamics, we refer the readers to Toba (1959); Princen (1963); Walls et al. (2015); Krishnan et al. (2017); Deike et al. (2018).
Figure 1: (a-i) A bubble with radius $R_{0}$ rests close to the liquid-gas interface, separated from it by a thin liquid film of thickness $\delta\ll R_{0}$. The surrounding viscoelastic medium is characterized by density $\rho_{s}$, solvent viscosity $\eta_{s}$, elastic modulus $G$, and relaxation time $\lambda$. The gas has density $\rho_{g}$ and viscosity $\eta_{g}$. (a-ii) Film rupture creates an axisymmetric cavity, which we study in this work. (b) Apart from the solvent Ohnesorge number $Oh_{s}=\eta_{s}/\sqrt{\rho_{s}\gamma R_{0}}$ and the Bond number $Bo=\rho_{s}gR_{0}^{2}/\gamma$, the presence of polymers introduces two additional parameters, namely the elastocapillary number $Ec=GR_{0}/\gamma$ (equation (3)) and the Deborah number $De=\lambda/\sqrt{\rho_{s}R_{0}^{3}/\gamma}$ (equation (4)). To explore the dynamics, we move across the entire $Ec$-$De$ phase space. Often, the polymeric Ohnesorge number $Oh_{p}=G\lambda/\sqrt{\rho_{s}\gamma R_{0}}{{\color[rgb]{0,0,0}=Ec\times De}}$ (equation (5)) based on polymeric viscosity is also used to describe the influence of polymers.
Given the potential for jet drops to transport pathogens or pollutants into the atmosphere, strategies to prevent their generation are pertinent. Recent studies unsurprisingly show that non-Newtonian effects, particularly viscoelasticity, can suppress jet drop production (Sanjay et al., 2021; Sen et al., 2021; Rodríguez-Díaz et al., 2023; Ji et al., 2023). While computational studies have successfully reproduced experimental observations, such as elasticity-induced droplet suppression (Cabalgante-Corralesa et al., 2024; Balasubramanian et al., 2024), the full impact of these effects on bubble-bursting dynamics remains elusive. Advancements in solving non-linear constitutive equations for highly deformed interfacial flows have been made possible by techniques like the log-conformation method (Fattal & Kupferman, 2004) and the square-root conformation method (Balci et al., 2011). Originally developed for single-phase flows, these methods have been extended to multiphase flows (Fraggedakis et al., 2016; López-Herrera et al., 2019; Varchanis & Tsamopoulos, 2022; França et al., 2024; Zinelis et al., 2023), facilitating more comprehensive investigations into this topic.
Viscoelastic media differ from viscous Newtonian liquids in their rheological properties, exhibiting both viscous and elastic stresses when deformed due to the presence of dissolved polymers. These polymeric effects are characterized by two material properties: the elastic modulus $G$ that characterizes the strength of the dissolved polymers by relating the strain with the additional polymeric stresses in the system, and the relaxation time scale $\lambda$ that characterizes the memory of the system as it is a measure of the timescale at which the additional polymeric stresses in the system vanish. When non-dimensionalizing these properties, we obtain two further non-dimensionalized control parameters, namely, the elastocapillary number
$Ec=\frac{GR_{0}}{\gamma},$ (3)
comparing the elastic modulus to the Laplace pressure scale, and the Deborah number
$De=\frac{\lambda}{\sqrt{\rho_{s}R_{0}^{3}/\gamma}},$ (4)
comparing the relaxation time of the additional stresses to the process timescale, i.e., the inertiocapillary timescale $\tau_{\gamma}=\sqrt{\rho_{s}R_{0}^{3}/\gamma}$. Additionally, we also introduce the polymeric viscosity $\eta_{p}=G\lambda$ based on dimensional arguments, which can be normalized with the inertiocapillary scales to give the polymeric Ohnesorge number (figure 1b)
$Oh_{p}=\frac{\eta_{p}}{\sqrt{\rho_{s}\gamma R_{0}}}={{\color[rgb]{0,0,0}Ec\times De}},$ (5)
which is the product of $Ec$ and $De$.
In this study, we investigate viscoelastic effects on bubble bursting dynamics by exploring the three-dimensional phase space of $Oh_{s}$, $Ec$, and $De$, using volume of fluid-based finite volume simulations. Our results demonstrate that the addition of polymers significantly influences the overall dynamics, which are governed by the interplay of viscous and elastic effects. For systems with a permanent memory of its initial state and subsequent deformations, i.e., when the additional polymeric stresses are sustained throughout the process time scale ($De\to\infty$), the dimensionless elastic modulus dictates the dynamics and suppression of jet and drops. In contrast, for systems with poor memory of its initial state and subsequent deformation ($De\to 0$), the dynamics resemble those encountered in Newtonian liquids with an effective viscosity deduced using the slender elastic jet equations.
Building upon the extensive literature on viscoelastic flows, we extend these concepts to the specific case of bubble bursting. Previous research has explored viscoelastic phenomena in various contexts, including flow through nozzles and contractions (Hinch, 1993; Chen, 1991; Boyko et al., 2024), stability and breakup of viscoelastic jets (Middleman, 1965; Goren & Gottlieb, 1982; Bousfield et al., 1986; Chang et al., 1999; Anna & McKinley, 2001; Pandey et al., 2021; Sen et al., 2024; Zinelis et al., 2023), coalescence and spreading of viscoelastic drops and bubbles (Bouillant et al., 2022; Dekker et al., 2022; Oratis et al., 2023), and oscillating bubbles in viscoelastic media (Oratis et al., 2024). Recent studies have also investigated elastoviscoplastic flows, incorporating viscous, elastic, and plastic aspects (Putz & Burghelea, 2009; Varchanis et al., 2019; França et al., 2024; Balasubramanian et al., 2024), further expanding our understanding of non-Newtonian liquids. We refer readers to reviews by Bogy (1979), Eggers (1997), and Yarin (1993) for comprehensive overviews of these topics. Our work applies the foundational knowledge developed in these works to elucidate how viscoelasticity alters the formation of Worthington jets and ejected droplets during bubble bursting, enhancing our understanding of this specific phenomenon.
This paper is organized as follows: § 2 presents the governing equations and numerical method. § 3 investigates the polymer influence on bubble bursting, focusing on systems with permanent ($De\to\infty$, § 3.1) and poor ($De\to 0$, § 3.2) memory. For both cases, we categorize bursting bubble dynamics into distinct regimes and elucidate the transitions in § 4 where we generalize the results across systems where the memory of the initial conditions and subsequent deformations is gradually fading ($0<De<\infty$). Finally, § 5 summarizes our findings and suggests future research directions.
2 Numerical framework and problem description
2.1 Governing Equation
We investigate the collapse of an open bubble cavity at the interface in a viscoelastic medium (of figure 1) using an axisymmetric domain with incompressible fluids. Length scales are normalized using the initial bubble radius giving =~$\mathcal{L}=\tilde{\mathcal{L}}R_{0}$ as characteristic length, and the time is normalized using the inertiocapillary timescale $\tau_{\gamma}=\sqrt{\rho_{s}{R_{0}}^{3}/\gamma}$ giving $t=\tilde{t}\tau_{\gamma}$. These normalizations yield an inertiocapillary velocity scale $u_{\gamma}=\sqrt{\gamma/\rho_{s}R_{0}}$ for the velocity field $\boldsymbol{u}=\tilde{\boldsymbol{u}}u_{\gamma}$. Lastly, all stresses are normalized using the Laplace pressure scale, $\boldsymbol{\sigma}=\tilde{\boldsymbol{\sigma}}\sigma_{\gamma}$, where $\sigma_{\gamma}=\gamma/R_{0}$. Here, as usual, non-dimensionalized quantities are denoted with a tilde, though from here onwards, we drop the tilde, and all equations are thus dimensionless in the current section. Throughout the manuscript, we use the subscripts $s$, $p$, and $g$ to denote liquid solvent, polymer, and gas, respectively. The governing mass and momentum conservation equations for the liquid phase read as
$\boldsymbol{\nabla\cdot u}=0,\,\text{and}$ (6)
[…]
where the Newtonian contribution (coming from the solvent) $\boldsymbol{\sigma_{s}}$ is
$\boldsymbol{\sigma_{s}}=2Oh_{s}\boldsymbol{\mathcal{D}},$ (8)
with $\boldsymbol{\mathcal{D}}=\left(\boldsymbol{\nabla u}+\left(\boldsymbol{\nabla u}\right)^{T}\right)/2$ representing the symmetric part of the velocity gradient tensor. The non-Newtonian contribution $\boldsymbol{\sigma_{p}}$ arises from the presence of polymers in the fluid. We emphasize that although we refer to $\boldsymbol{\sigma_{p}}$ as ‘polymeric stresses’ in the context of dilute polymer liquids, this concept extends to any deformable microstructure within the fluid that responds to flow (Saramito, 2007; Snoeijer et al., 2020; França et al., 2024; Balasubramanian et al., 2024). To characterize the deformation of these microstructures, we introduce the conformation tensor $\boldsymbol{\mathcal{A}}$, an order parameter that evolves from an initial identity state $\boldsymbol{\mathcal{A}}=\boldsymbol{\mathcal{I}}$ (figure 1a-ii). For the stress-strain constitutive relationship, we employ the model based on the Oldroyd-B family of models, which represents the simplest conformation tensor-based approach for viscoelastic fluids (Oldroyd, 1950; Bird et al., 1977; Snoeijer et al., 2020; Stone et al., 2023; Boyko & Stone, 2024). This model assumes a linear relationship between elastic stresses and polymeric deformation,
$\boldsymbol{\sigma_{p}}=Ec\left(\boldsymbol{\mathcal{A}}-\boldsymbol{\mathcal{I}}\right),$ (9)
where $Ec$ is the elastocapillary number (equation (3)), representing the strength of the polymers analogous to a dimensionless elastic modulus. Note that even though the polymeric stresses $\boldsymbol{\sigma_{p}}$ grow linearly with $\boldsymbol{\mathcal{A}}$, the polymeric deformations $\boldsymbol{\mathcal{A}}$ can be highly nonlinear. Naturally, in the limit of $Ec=0$, the polymeric stress would vanish, and the system will give a viscous Newtonian dictated by the solvent Ohnesorge number $Oh_{s}$ (see equation (8)).
Additionally, the conformation tensor $\boldsymbol{\mathcal{A}}$ relaxes to its base state $\boldsymbol{\mathcal{I}}$ over time due to thermal effects. Once more, using the Oldroyd-B model, $\boldsymbol{\mathcal{A}}$ follows a linear relaxation law,
[…]
where
${{\boldsymbol{\mathcal{A}}}}\equiv\frac{\partial\boldsymbol{\mathcal{A}}}{\partial t}+\left(\boldsymbol{u\cdot\nabla}\right)\boldsymbol{\mathcal{A}}-2\text{Sym}\left(\boldsymbol{\mathcal{A}\cdot}\left(\boldsymbol{\nabla u}\right)\right)$ (11)
is the frame-invariant upper convected derivative of second-rank tensor $\boldsymbol{\mathcal{A}}$, and $De=\lambda/\tau_{\gamma}$ (defined in equation (4)) is the Deborah number, representing the ratio of the polymer relaxation time $\lambda$ to the process timescale $\tau_{\gamma}$ and Sym depicts the symmetric part of the tensor. The Deborah number characterizes the polymeric liquid’s memory. It is instructive to note that in the limit of $De\to\infty$, polymeric liquids have permanent memory and the dissolved polymers undergo affine motion (see equation (10) and Snoeijer et al., 2020; Stone et al., 2023; Boyko & Stone, 2024)
[…]
indicating that they follow the flow and deform according to the velocity field. In this limit, for finite $Ec$, the Oldroyd-B model is equivalent to the damped neo-Hookean model (also known as the Kelvin-Voigt model) for solids (Snoeijer et al., 2020). Conversely, at $De=0$, polymeric liquids have no memory of their initial condition and subsequent deformations, relaxing immediately to the base state. For non-infinite $Ec$ values, polymeric stresses vanish, resulting in a Newtonian response (equation (9)) governed by the solvent Ohnesorge number $Oh_{s}$ (see equation (8)). It is, therefore, surprising that both $Ec=0$ and $De=0$ (figure 1b) represent Newtonian responses, irrespectively of the corresponding other parameter.
Equations (9) and (10) can be combined to get
[…]
where $Oh_{p}=Ec\times De$ is the polymeric Ohnesorge number (equation (5)). Consequently, in the limit $De\to 0$ at fixed $Oh_{p}$ (e.g., moving along constant $Oh_{p}$ lines in figure 1b), the system exhibits a viscous Newtonian response with a total dimensionless viscosity of $Oh_{s}+Oh_{p}$.
The Oldroyd-B model, despite its widespread use due to its simplicity, fails to capture several important physical phenomena (Snoeijer et al., 2020). It inadequately describes shear-thinning behavior in polymeric liquids (Yamani & McKinley, 2023) and erroneously predicts unbounded stress growth in strong extensional flows (McKinley & Sridhar, 2002; Eggers et al., 2020). These limitations can be addressed by incorporating finite polymer extension, for example, by increasing the effective $Ec$ as the polymer approaches full extension (Hinch & Harlen, 2021; Zinelis et al., 2023). Various extensions of the Oldroyd-B equations have been developed to account for such nonlinearity, either in equations (9) and (10) or in the solvent contribution in equation (8) (Tanner, 2000). In this study, we employ the Oldroyd-B model to include the two primary effects of the polymer addition: the additional stress ($Ec$) and polymeric liquid memory ($De$). Our aim is to provide a comprehensive understanding of the entire $Ec$-$De$ parameter space (figure 1b).
2.2 Methods
We employ the open-source software Basilisk C (Popinet & collaborators, 2013–2024; Popinet, 2015) to solve the governing equations outlined in § 2.1. The equations are solved using the one-fluid approximation (Tryggvason et al., 2011), with surface tension incorporated as singular body force at the liquid-gas interface (Brackbill et al., 1992). To account for the gas phase, in addition to the dimensionless parameters described in § 1 and § 2.1, we maintain constant density and viscosity ratios of $\rho_{r}=\rho_{g}/\rho_{s}=10^{-3}$ and $\eta_{r}=\eta_{g}/\eta_{s}=2\times 10^{-2}$, respectively. The liquid-gas interface is tracked using the volume of fluid (VoF) method, governed by the advection equation
$\frac{\partial\Psi}{\partial t}+\boldsymbol{\nabla\cdot}\left(\Psi\boldsymbol{u}\right)=0,$ (14)
where $\Psi$ represents the VoF color function. We implement a geometric VoF approach, reconstructing the interface at each timestep and applying surface tension forces as singular forces (Popinet, 2009; Brackbill et al., 1992)
$\boldsymbol{f}{\gamma}\approx\kappa\boldsymbol{\nabla}\Psi,$ (15)
with curvature $\kappa$ calculated using the height-function method (Popinet, 2018). The explicit treatment of surface tension imposes a time step constraint based on the smallest capillary wave oscillation period (Popinet, 2009). Yet another time step restriction, usually more relaxed than the surface tension one, comes from the explicit treatment of the polymeric stress term $\boldsymbol{\sigma}_{p}$. We impose no-penetration and free-slip conditions at wall boundaries to avoid wall-shear effects, with outflow conditions at the top boundary to prevent droplet rebound. Pressure gradients are set to zero at domain boundaries for both liquid and gas phases.
The initial bubble shape is determined by solving the Young-Laplace equations for quasi-static equilibrium (Princen, 1963; Toba, 1959; Villermaux et al., 2022; Sanjay, 2022). While the shape’s asymmetry increases with the Bond number $Bo$, we focus on the limit $Bo\to 0$, setting $Bo=0.001$ to regularize the singularity at the sphere-plane intersection. This results in a near-spherical initial cavity shape (figure 1a-i). During the bubble cap bursting, the film cap retracts almost instantaneously, after which the capillary waves are generated. As we are interested only in the bubbe cavity collapse, the simulations begin with an open cavity without the thin cap (figure 1a-ii), as also done similarly in recent studies (Deike et al., 2018; Gordillo & Rodríguez-Rodríguez, 2019; Sanjay et al., 2022). The computational domain spans $8R_{0}\times 8R_{0}$, discretized using quadtree grids with adaptive mesh refinement (AMR) (Popinet, 2009). Error tolerances for the VoF color function, curvature, velocity, and order parameter $\boldsymbol{\mathcal{A}}$ are set to $10^{-3}$, $10^{-6}$, $10^{-3}$, and $10^{-3}$, respectively.
In this work, following our earlier study (Sanjay et al., 2022), most simulations maintain a minimum grid size of $\Delta=R_{0}/512$, which dictates that, to get consistent results, 512 cells are required across the bubble radius while using uniform grids. We have also used an increased resolution ($\Delta=R_{0}/1024$ for high $De$ cases and $\Delta=R_{0}/2048$ near transitions) as needed. We have carried out extensive grid independence studies to ensure that changing the grid size does not influence the results. We refer the readers to Popinet (2015); Sanjay (2022); Sanjay & Dixit (2024) for further details of the numerical method used in this work.
Figure 2: Temporal evolution of the bubble cavity collapse at $De\to\infty$ and $Oh_{s}=0.025$ for $Ec=$ (a) $0.0001$, (b) $0.01$, and (c) $0.1$. The color scheme in the left panel of each snapshot represents the magnitude of the velocity field normalized by the inertiocapillary velocity, while on the right panel of each snapshot, it shows the trace of the elastic stress $\boldsymbol{\sigma}_{p}$ that represents twice the elastic energy stored in polymeric deformations on a $\log_{10}$ scale. See also the supplementary movies SM1.
3 Influence of polymers
This section phenomenologically describes the influence of polymers on the bursting bubble process by investigating how varying the elastocapillary number $Ec$ influences the formation of Worthington jets and droplet ejection. We focus on two limiting cases: polymeric solutions with permanent memory exhibiting affine motion ($De\to\infty$) and those with poor memory ($De\to 0$).
3.1 Polymeric liquids with permanent memory
We begin our analysis by considering the limit of $De\to\infty$, where the polymeric solutions feature affine motion (equation (12)) and maintain a permanent memory of their initial condition and subsequent deformations without relaxation during the process timescale. Figure 2 illustrates representative cases in viscoelastic media for $Oh_{s}=0.025$ and varying elastocapillary numbers ($Ec$). The figure presents a temporal evolution of the interface profile (green line) alongside with the velocity magnitude on the left and the trace of elastic stress $\boldsymbol{\sigma}_{p}$ on the right. Remarkably, despite all cases exhibiting a total Ohnesorge number of infinity ($Oh_{s}+Oh_{p}\to\infty$), which typically implies highly viscous behavior (see figure A 1), low $Ec$ scenarios demonstrate dynamics qualitatively resembling Newtonian fluids. In these cases, capillary waves drive the collapse of a bubble cavity, converging at its bottom to form a Worthington jet that subsequently fragments into droplets (see figure 2a). Intuitively, the elastic stresses are concentrated near the axis of symmetry where the strain is maximum (Turkoz et al., 2018; Eggers et al., 2020). The process concludes within a finite timescale ($\sim\tau_{\gamma}$), resulting in a regular limit as $Ec\to 0$. As a result, the system’s behavior deviates gradually from the Newtonian case at $Ec=0$, exhibiting a continuous transition as the elasticity increases. This absence of singularity contrasts with elastic Taylor–Culick-type retractions, where an infinite process timescale allows the elastic stresses to develop, leading to distinct behaviors for $Ec=0$ and $Ec\to 0$ (Bertin et al., 2024), i.e., a singular limit.
We stress that in this limit, the jet breakup occurs due to finite grid resolution in our numerical code (Lohse & Villermaux, 2020; Chirco et al., 2022; Kant et al., 2023). We cannot differentiate between a case of drop detachment from the jet or the case when they are still connected through a thin filament–also known as the beads-on-a-string structure (Hosokawa et al., 2023; Clasen et al., 2006; Pandey et al., 2021; Zinelis et al., 2023). Although current simulations fully resolve other aspects, they cannot resolve these finest threads, which may have subgrid cell sizes depending on the $Ec$. At higher grid resolutions, we expect to recover the beads-on-a-string configuration, as the Oldroyd-B model does not yield a finite time breakup singularity in the infinite $De$ regime, instead converging to a finite filament (Eggers et al., 2020; Turkoz et al., 2018, 2021). To prevent infinite thread thinning, a nonlinear elastic model could also be employed (see § 2.1 for further discussions).
As $Ec$ increases, we observe jet formation without droplet ejection (figure 2b). At higher $Ec$ values, even jet formation is suppressed due to elevated elastic resistance (figure 2c). Notably, while polymeric effects significantly influence the dynamics after the convergence of capillary waves (figure 2, $t/\tau_{\gamma}=0.8,1.2$), the propagation of capillary waves (figure 2, $t/\tau_{\gamma}=0.1,0.4$) remains largely unaffected. Figure 3(a) quantifies the trajectories of these capillary waves across three orders of magnitude variation in $Ec$ at two different $Oh_{s}$. The capillary wave speed is independent of both liquid and polymeric control parameters, mirroring the behavior observed in Newtonian media (Gordillo & Rodríguez-Rodríguez, 2019) and contrasting those for viscoplastic media (Sanjay et al., 2022). Following capillary wave collapse, the Worthington jet initially elongates to a maximum length ($L_{\text{max}}$) before retracting. As shown in Figure 3(b) for $Oh_{s}=0.04$, $L_{\text{max}}$ decreases with increasing $Ec$ due to stronger resistive stresses.
Figure 3: (a) Trajectory of the maximum curvature capillary wave parameterized using the angle $\theta_{c}(t)$ as depicted in the inset at $De\to\infty$ for different $Oh_{s}$ and $Ec$. (b) Evolution of the jet length $L(t)$ at $Oh_{s}=0.04$ and $De\to\infty$ for different $Ec$.
Figure 4(a) presents a phase map of $L_{\text{max}}$, compiled from approximately 100 simulations. For Newtonian liquids, $L_{\text{max}}$ peaks near $Oh_{s}\approx 0.03$, corresponding to the value of observed hydrodynamic singularities (Zeff et al., 2000; Lohse, 2003; Eggers & Fontelos, 2015; Yang et al., 2020), before decreasing at higher $Oh_{s}$ (Duchemin et al., 2002; Deike et al., 2018; Gordillo & Rodríguez-Rodríguez, 2019). Jet formation ceases altogether beyond a critical value of $Oh_{c}=0.11$ (Sanjay et al., 2021) (defined here when $L_{\text{max}}<0.3R_{0}$). As $Ec$ increases, viscoelastic effects become significant. $L_{\text{max}}$ decreases monotonically with $Ec$ due to increased elastic resistance, with jet formation suppressed beyond $Ec=0.086$. Unlike the non-monotonic relationship between $L_{\text{max}}$ and $Oh_{s}$, where increasing $Oh_{s}$ initially produces thinner and faster jets, the $L_{\text{max}}(Ec)$ relationship remains consistently monotonic. Even the $Oh_{s}$-sensitive singular Worthington jets disappear with increasing $Ec$. Notably, the critical $Ec$ values for these transitions appear to be largely independent of $Oh_{s}$, in contrast to the $Oh_{s}$-dependent behavior observed in the Newtonian limit.
Figure 4: (a) The maximum jet length $L_{\text{max}}$ at $De\to\infty$ in the $Ec$-$Oh_{s}$ phase space, depicted by the colormap, where the lighter region corresponds to higher values. For the Newtonian liquid $\left(Ec\to 0\right)$, the jetting transition occurs at $Oh_{s}=0.11$, denoted by the horizontal dotted line. Due to the elastic effects, this transition occurs at $Ec=0.086$, as depicted by the vertical dotted line. (b) The size of the first droplet at $De\to\infty$ in the $Ec$-$Oh_{s}$ phase space. For the Newtonian liquid, the dropping transition is observed at $Oh_{s}=0.0375$, denoted by the horizontal dotted line. Further, the transition due to elastic effects is very sensitive to $Oh_{s}$ and is shown by the inclined dotted line.
The emerging Worthington jet may eject multiple droplets. For Newtonian liquids, exact solutions for the first droplets’ size $r_{d}$ are well understood (see appendix A and Blanco-Rodríguez & Gordillo (2020)). $r_{d}$ decreases with $Oh_{s}$ until $Oh_{s}\approx 0.0375$, beyond which the droplet breaks from the jet due to the Rayleigh–Plateau instability and falls downwards. Our analysis focuses on droplets propagating away from the source, excluding those with downward velocity upon breakup (observed in Newtonian media for $0.0450.0375<Oh_{s}<$ 0.045). For elastic cases, despite unresolved filaments connecting droplets and jets, we have rigorously verified the convergence of the first droplet’s size to at least 10% accuracy. Figure 4(b) illustrates a phase map of the first droplet’s size $r_{d}$, revealing intriguing differences from the jet behavior. While $r_{d}$ follows the same trend with $Oh_{s}$ and remains invariant of $Ec$ below critical values, the critical $Ec$ for droplet suppression differs from that of jet suppression. As the jet width is determined solely by $Oh_{s}$, independently of $Ec$, the first emerging droplet’s size also remains independent initially. However, as $Ec$ increases further, rising elastic stresses suppress droplet formation more abruptly than jet formation. The critical values $Ec_{d}$ for the transition between jet formation with and without droplet breakup (dropping transition) are sensitive to $Oh_{s}$, with the critical $Ec_{d}$ decreasing as $Oh_{s}$ increases. This trend is in stark contrast with the transition from jet formation to jet suppression (jetting transition), where critical $Ec$ values remain largely $Oh_{s}$-independent.
Figure 5: Temporal evolution of bubble cavity collapse at $De=0.01$ and $Oh_{s}=0.025$ for $Ec=$ (a) $1$, (b) $5$, and (c) $10$. The color scheme in the left panel of each snapshot represents the magnitude of the velocity field normalized by the inertiocapillary velocity, while on the right panel of each snapshot, it shows the trace of the elastic stress $\boldsymbol{\sigma}_{p}$ that represents twice the elastic energy stored in polymeric deformations on a $\log_{10}$ scale. See also the supplementary movies SM2.
Figure 6: (a) The maximum jet length $L_{\text{max}}$ at $De=0.01$ in the $Ec$-$Oh_{s}$ phase space, depicted by the colormap, where the lighter region corresponds to higher values. For the Newtonian liquid, the jetting transition occurs at $Oh_{s}=0.11$, denoted by the horizontal dotted line. Due to the elastic effects, this transition occurs at $Ec=9.3$, as depicted by the vertical dotted line. (b) The size of the first droplet at $De=0.01$ in the $Ec$-$Oh_{s}$ phase space. For the Newtonian liquid, the dropping transition is observed at $Oh_{s}=0.0375$, denoted by the horizontal dotted line. Further, the $Oh_{s}$-independent transition due to elastic effects occurs at $Ec=2.5$, as shown by the vertical dotted line.
3.2 Polymeric liquids with poor memory
This section examines the dynamics in media with a poor memory of its initial conditions and subsequent deformations. ($De\to 0$). For sufficiently small Deborah numbers $De$, the polymers relax rapidly, resulting in elastic stresses of the polymeric liquid that are considerably lower than those observed in cases where $De\to\infty$. The stress relaxation also results in the dissipation of elastic energy stored in stretched polymers. Figure 5 illustrates representative cases for $De=0.01$, showcasing three distinct regimes as a function of the elastocapillary number ($Ec$). The figure presents a temporal evolution of the interface profile (green line) alongside velocity magnitude on the left and the trace of elastic stress $\boldsymbol{\sigma}_{p}$ on the right for $Ec=1$, $5$, and $10$. For $Ec=1$ (figure 5a), we observe a slender Worthington jet that forms a droplet. As $Ec$ increases to $5$ (figure 5b), the jet persists but fails to produce a droplet. At $Ec=10$ (figure 5c), jet formation is completely suppressed, with the interface showing only slight deformations during cavity relaxation. The qualitative trends with respect to the elastocapillary number ($Ec$) remain consistent as compared to those in § 3.1. However, the critical $Ec$ values for different regimes differ markedly from those observed at $De\to\infty$. Notably, jet formation and droplet production persist at $Ec=1$ (figure 5a), despite this value being an order of magnitude higher than the critical $Ec$ for the jetting transition at infinite $De$. This difference underscores the dependence of transition thresholds on $De$.
To further interpret the jetting dynamics and drop formation, figure 6 presents phase maps illustrating the behavior of maximum jet lengths ($L_{\text{max}}$) and first droplet sizes ($r_{d}$) for $De=0.01$. Figure 6(a) shows $L_{\text{max}}$ across a range of $Ec$ and $Oh_{s}$ values. For low $Ec$, $L_{\text{max}}$ shows Newtonian-like $Oh_{s}$ dependence. As $Ec$ increases, $L_{\text{max}}$ decreases monotonically until jet formation ceases beyond an $Oh_{s}$-independent critical $Ec_{j}$, mirroring the infinite $De$ limit behavior. Figure 6(b) maps the $r_{d}$, showing $Ec$-independent droplet sizes until near the transition point, where droplet formation is suppressed. For $De\ll 1$, the critical $Ec_{d}$ for the dropping transition exhibits minimal $Oh_{s}$-dependence, contrasting with the $Oh_{s}$-sensitive behavior at infinite $De$. Comparing these results to the $De\to\infty$ limit reveals persistent fundamental regimes across different $De$ values, but the transition thresholds are highly sensitive to the polymeric liquid’s relaxation time. Critical $Ec$ values for both jet and droplet suppression are significantly higher at low $De$ compared to the infinite $De$ limit, indicating that rapid relaxation of polymeric stresses allows jet and droplet formation at higher $Ec$ values. This low $De$ behavior suggests an interplay between elastic and viscous effects, explored further in § 4.
4 Regime Map
The bursting bubble dynamics in viscoelastic media exhibit distinct behavior compared to Newtonian fluids. While capillary wave propagation remains largely unaffected by viscoelasticity, the Worthington jet dynamics are highly sensitive to both the elastocapillary number $Ec$ and Deborah number $De$. This sensitivity manifests in three distinct regimes: (i) jets forming droplets, (ii) jets without droplet formation, and (iii) absence of jet formation. By extending our analysis across the entire $Ec$-$De$ phase space, we elucidate the transitions between these regimes, building upon our previous examination of the limiting cases $De\to\infty$ and $De\to 0$ from § 3.
4.1 Summary of the different regimes
Figure 7: The elastocapillary-Deborah number ($Ec$-$De$) phase map delineating the transition between the regimes: (i) jets forming droplets and (ii) jets without droplet formation. The data points represent the critical elastocapillary number $Ec_{d}(De,Oh_{s})$ at which this transition occurs. The transition behavior exhibits distinct characteristics in different limits: as $De\to\infty$, the transition occurs at a constant $Ec$ which is highly sensitive to $Oh_{s}$ (see the gray dashed line showing $Ec_{d}\sim De^{0}$), while for $De\to 0$, the transition is $Oh_{s}$-independent and occurs at constant $Oh_{p}$ (see the gray solid line showing $Ec_{d}\sim De^{-1}$, i.e., $Oh_{p,d}\sim De^{0}$).
Figure 8: (a) The elastocapillary-Deborah number ($Ec$-$De$) and (b) the polymeric Ohnesorge-Deborah number ($Oh_{p}$-$De$) phase map delineating the transition between the regimes: (ii) jets without droplet formation and (iii) absence of jet formation. The data points represent the $Oh_{s}$-independent critical elastocapillary number $Ec_{j}(De)$ at which this transition occurs. The transition behavior exhibits distinct characteristics in different limits: as $De\to\infty$, the transition occurs at a constant $Ec$ (see gray dashed line showing $Ec_{d}\sim De^{0}$), while for $De\to 0$, the transition occurs at constant $Oh_{p}$ (see gray solid line showing $Ec_{d}\sim De^{-1}$, i.e., $Oh_{p,d}\sim De^{0}$).
Figure 7 depicts the elastocapillary-Deborah number ($Ec$-$De$) phase diagram, distinguishing between jets forming droplets and those without droplet formation. Figure 8 further delineates the transition to the absence of jet formation. These transitions exhibit distinct characteristics at two asymptotic limits: $De\to\infty$ and $De\to 0$. For polymeric liquids with long relaxation times ($De\gg 1$):
  1. 1.
    the dropping transition occurs at $Ec_{d}(Oh_{s})$, exhibiting strong dependence on $Oh_{s}$ (Figures 4b and 7), and
  2. 2.
    the jetting transition takes place at $Ec_{j}\approx 0.086$, independently of $Oh_{s}$ (Figure 8a).
Notably, both transitions at $De>1$ occur asymptotically close to their corresponding solutions at $De\to\infty$, reflecting the finite process duration. When the relaxation time exceeds the process timescale, polymers lack sufficient time to relax, rendering the process $De$-independent.
For polymeric liquids with short relaxation times ($De\ll 1$), both transitions are $Oh_{s}$-independent and occur at constant polymeric Ohnesorge number $Oh_{p}=Ec\times De$:
  1. 1.
    the dropping transition occurs at $Oh_{p,d}\approx 0.048$ (Figure 7) and
  2. 2.
    the jetting transition occurs at $Oh_{p,j}\approx 0.129$ (Figure 8b).
These distinct transition behaviors reveal that for very small $De$, the addition of polymers can be considered to have an added viscous effect characterized by the polymeric Ohnesorge number $Oh_{p}$. In contrast, at high $De$, the medium behaves analogously to an elastic solid for $Oh_{s}\to 0$ and as a Kelvin-Voigt solid for finite $Oh_{s}$. We further investigate the jetting transition using slender jet equations in § 4.2 following similar approaches by Driessen et al. (2013); Gordillo et al. (2020); Zinelis et al. (2023); Sen et al. (2024).
Figure 9: Temporal evolution of the Worthington jet for a representative case, where the jet emerges, reaches a maximum, and is pulled to merge with the liquid bath. The control volume contains the jet region, as shown by the region within the gray lines. Here, $h(z,t)$ is the width of the jet, which becomes $h_{\text{base}}$ at the base of the jet. The capillary force at the jet base is $F_{\gamma}=\gamma\left(2\pi h_{\text{base}}\right)$ that acts radially outwards. At the same time, the elastic and viscous stresses act at the base of the jet as ${{\color[rgb]{0,0,0}F_{\eta}+F_{p}=\left(\sigma_{\eta,\text{base}}+\sigma_{p,\text{base}}\right)\pi h_{\text{base}}^{2}}}$.
4.2 What sets the different transitions, and what do we learn from these transitions?
To understand the mechanisms governing bubble cavity collapse, we analyze jet dynamics using a control volume approach (figure 9). Employing the slender jet approximation (Eggers & Fontelos, 2015, p. 240), given the small radial-to-axial length scale ratio, the vertical momentum equation for the jet reads
$\rho_{s}\left(\frac{\partial v}{\partial t}+v\frac{\partial v}{\partial z}\right)=-\gamma\frac{\partial\kappa}{\partial z}+\frac{1}{h^{2}}\frac{\partial}{\partial z}\left[h^{2}\left(3\eta_{s}\frac{\partial v}{\partial z}+G\left(\mathcal{A}_{zz}-1\right)\right)\right].$ (16)
Here, $v(z,t)$ is the radially averaged jet velocity, and the shape of this jet is $h(z,t)$. We define a control volume containing the emerging jet that is always bounded by the inflection points at the interfaces, see figure 9b. Integrating over this control volume (with differential volume element $d\Omega=\pi h(z,t)^{2}\mathrm{d}z$) yields the force balance (Trouton, 1906):
$\frac{d\mathcal{M}_{\text{jet}}}{dt}=3\eta_{s}h^{2}\frac{\partial v}{\partial z}\Bigg{|}_{\text{base}}+Gh^{2}(\mathcal{A}_{zz}-1)\Bigg{|}_{\text{base}}={{\color[rgb]{0,0,0}\left(\sigma_{\eta,\text{base}}+\sigma_{p,\text{base}}\right)\pi h_{\text{base}}^{2}}}$ (17)
where $\mathcal{M}_{\text{jet}}(t)=\int_{\Omega(t)}\rho_{s}v(z,t)\pi h(z,t)^{2}\mathrm{d}z$ denotes the momentum of the jet. The capillary stress (first term on the right-hand side of equation (16)) integral vanishes due to orthogonal interface intersection with the control volume (see Marchand et al. (2011) and p. 16-21, Munro (2019)). We chose this control volume because of its vanishing integral feature. Furthermore, the integral of the second term on the right-hand side forms an exact integral which vanishes at the tip where it is zero owing to $h(z=L_{\text{max}}(t))=0$. Consequently, jet evolution depends solely on stresses at the base: viscous ($\sigma_{\eta,\text{base}}(t)$) and elastic ($\sigma_{p,\text{base}}(t)$). For relevant $Oh_{s}$ values, $\sigma_{\eta,\text{base}}(t)$ is too weak to suppress the Worthington jet. Numerical simulations allow us to estimate $\sigma_{p,\text{base}}(t)$. As the capillary waves collapse, the base elastic stress reaches a global maximum, before decreasing again at later times. Jet formation occurs if inertial flow focusing is sufficiently strong at the peak elastic stress. We will now evaluate this competition for the two limits of $De$.
[…]
4.2.1 The limit of $De\to\infty$
Figure 10(a) shows that for $De>1$, the maximum elastic stress $\text{max}\left(\sigma_{p,\text{base}}(t)\right)$ reaches a plateau, dependent only on $Ec$. This $De$-independence coincides with the extent of infinite $De$ asymptotes featured in the transitions discussed in § 4.1. The upper limit of elastic resistance competes with inertial flow focusing to inhibit jet formation. We quantify the inertial stresses at peak elastic stress using:
$\sigma_{I,\text{base}}=\frac{2}{h_{\text{base}}^{2}}\int_{o}^{h_{\text{base}}}\rho_{s}v^{2}h\mathrm{d}h,$ (18)
where $h_{\text{base}}$ is the jet width at its base (see figure 9). Figure 10(b) reveals that the ratio of elastic to inertial stresses is largely independent of $Oh_{s}$. As $Ec$ increases, this ratio reaches a maximum beyond which jet suppression occurs. It is important to note that the apparent decrease in this stress ratio with increasing $Ec$ and $Oh_{s}$ near the jetting transition in figure 10(b) occurs due to a decrease in both the polymeric and inertial stresses in this region of the parameter space.
Figure 11: (a) Evolution of the maximum elastic stress at the jet base ($\text{max}\left(\sigma_{p,\text{base}}(t)\right)$), normalized by the Newtonian-like viscous stress $\sigma_{N,\text{base}}$ with viscosity $\eta_{p}=G\lambda$, as a function of $De$ for different $Oh_{p}$ at $Oh_{s}=0.001$. The gray dashed horizontal line represents $\text{max}\left(\sigma_{p,\text{base}}(t)\right)\approx\sigma_{N,\text{base}}$ while the black dashed line serves as a guide to the eye representing $\text{max}\left(\sigma_{p,\text{base}}(t)\right)/\sigma_{N,\text{base}}\sim De$. (b) The variation of jet’s tip velocity $V_{\text{jet}}$, normalized by the inertiocapillary velocity $u_{\gamma}=\sqrt{\gamma/\rho_{s}R_{0}}$, with $Oh_{\text{effective}}=3Oh_{s}+2Oh_{p}$ at different $De$ and $Oh_{s}=0.01$. The gray dashed line represents $V_{\text{jet}}\sim\gamma/\eta_{\text{effective}}$.
4.2.2 The limit of $De\to 0$
In the limit of $De\to 0$, polymeric liquids exhibit an added viscous effect characterized by the polymeric Ohnesorge number $Oh_{p}$ (also see § 4.1). To quantify this stress, we evaluate the maximum elastic stress $\text{max}(\sigma_{p,\text{base}}(t))$ normalized by a Newtonian-like viscous stress $\sigma_{N,\text{base}}$, calculated when the elastic stress reaches its maximum using
$\sigma_{N,\text{base}}=\frac{2}{h_{\text{base}}^{2}}\int_{o}^{h_{\text{base}}}G\lambda\frac{\partial v}{\partial z}h\mathrm{d}h,$ (19)
where $\eta_{p}=G\lambda$ represents the polymeric contribution to the shear viscosity of the polymeric liquid. Figure 11(a) demonstrates that $\text{max}(\sigma_{p,\text{base}}(t))$ collapse for all $Oh_{p}$ when rescaled with $\sigma_{N,\text{base}}$ for $De\to 0$.
As $De$ approaches unity, marking the onset of the infinite $De$ asymptotic regime, we observe a scaling behavior of $\text{max}(\sigma_{p,\text{base}}(t))\sim De\times\sigma_{N,\text{base}}$. This scaling remarkably resembles that predicted by Boyko et al. (2024) for flow in a slowly varying contraction at the infinite $De$ asymptote, despite significant geometric differences. While our study focuses on free surface flows and Boyko et al. (2024) examined contraction geometries, this unexpected similarity hints at a potentially universal behavior near the infinite $De$ asymptote. To further examine this intriguing connection, a similar closed-form $De$ expansion for free surface flows is necessary. However, we caution that this scaling approach to the infinite $De$ asymptote could be system-dependent (Hinch et al., 2024).
Here, we focus at the zero $De$ limit, where the elastic stress can be replaced with a Newtonian-like viscous stress with polymeric viscosity $\eta_{p}$ such that the jetting transition occurs at constant polymeric $Oh_{p}$. The polymeric stress contribution becomes $\sigma_{p}\approx 2G\lambda\boldsymbol{\mathcal{D}}$ (for the Oldroyd-B rheology, see also equation (13)). Consequently, we can rewrite the force balance in equation (17) as
$\frac{d{{\color[rgb]{0,0,0}\mathcal{M}}}_{\text{jet}}}{dt}=\left(3\eta_{s}+2G\lambda\right)h^{2}\frac{\partial v}{\partial z}\Bigg{|}_{\text{base}}{{\color[rgb]{0,0,0},}}$ (20)
which depicts the balance of jet inertia with viscous forces. Using as relevant scales for the jet momentum $\mathcal{M}_{\text{jet}}\sim\rho_{s}V_{\text{jet}}h_{\text{base}}^{3}$, where $V_{\text{jet}}$ is jet velocity, the velocity gradient $\partial_{z}v\sim V_{\text{jet}}/\delta_{\eta}$, and the time scale $\tau_{i}\sim h_{\text{base}}/V_{\text{jet}}$, we obtain the balance of forces as
$\rho V_{\text{jet}}^{2}\sim\eta_{\text{effective}}\frac{V_{\text{jet}}}{\delta_{\eta}}{{\color[rgb]{0,0,0}.}}$ (21)
Here, $\delta_{\eta}$ is the relevant length scale for viscous forces, and
$\eta_{\text{effective}}=3\eta_{s}+2G\lambda$ (22)
is the effective viscosity. As polymers do not affect the flow before jet formation (§ 3), the jet Weber number remains constant at inception (Gordillo & Rodríguez-Rodríguez, 2019),
$We_{\text{jet}}=\frac{\rho V_{\text{jet}}^{2}\delta_{\eta}}{\gamma}=\,\text{constant}.$ (23)
Combining equations (21) and (23), we get
$V_{\text{jet}}\sim\frac{\gamma}{\eta_{\text{effective}}}$ (24)
which has the same form as for Newtonian media (Gordillo & Rodríguez-Rodríguez, 2019), albeit with a modified viscosity term. This scaling holds in the viscous regime for Newtonian liquids (Blanco-Rodríguez & Gordillo, 2020).
Figure 12: The capillary waves focus and collapse at the bottom of the cavity. (a) The inception of the jet after the collapse at different $Oh_{p}$ at $De=0.001$ and $Oh_{s}=0.01$. The radius of the jet at the base $R_{\text{jet}}$ decreases with $Oh_{p}$ until $Oh_{p,c}=0.017$, beyond which bubbles are entrained. (b) Radius of jet $R_{\text{jet}}$ with $Oh_{p}$ at $De=0.001$ and $Oh_{s}=0.01$. $R_{\text{jet}}$ remains close to the value at the Newtonian limit $Oh_{p}=0$, and decreases sharply as it approaches $Oh_{p,c}$. Beyond $Oh_{p,j}$ jets are no longer observed.
Figure 11(b) illustrates the jet velocity as a function of the effective Ohnesorge number
$Oh_{\text{effective}}=3Oh_{s}+2Oh_{p}$ (25)
(reflecting equation (22)) at different $De$. We stress that the jet velocity varies in time (Deike et al., 2018; Sanjay et al., 2022; Gordillo & Blanco-Rodríguez, 2023) and is maximum at its inception, which is the value that we report here. For sufficiently large $Oh_{\text{effective}}$ and small $De$, we recover the scaling predicted in equation (24). However, as $De$ increases, the added elastic stresses cannot be directly substituted with Newtonian-like viscous stresses, and the underlying assumption fails, evident in the deviation of $V_{\text{jet}}$ from the prediction.
On the other hand, for small $Oh_{\text{effective}}$, $V_{\text{jet}}$ for all $De$ closely matches the corresponding speed in Newtonian liquids, as observed in figure 11b for $Oh_{p}=0$. As $Oh_{p}$ increases, $V_{\text{jet}}$ also increases, reaching a maximum before decreasing and following equation (24). Although the capillary wave speed remains unaffected in the polymeric medium, increasing $Oh_{p}$ triggers elastic stresses in smaller wavelength capillary waves, which are promptly dissipated due to small $De$. Consequently, improved flow focusing occurs as the strongest undamped capillary wave survives, thus increasing $V_{\text{jet}}$. This behavior is analogous to the non-monotonicity observed and well-understood for Newtonian liquids at small $Oh_{s}$ (Duchemin et al., 2002; Deike et al., 2018; Gordillo & Rodríguez-Rodríguez, 2019; Sanjay et al., 2022; Yang et al., 2020; Gordillo & Blanco-Rodríguez, 2023), further supporting the observation that polymeric liquid exhibit a Newtonian-like viscous response in the zero $De$ limit.
To further quantify this behavior, figure 12(a) illustrates jet features at inception for different $Oh_{p}$, while figure 12(b) shows jet radius as a function of $Oh_{p}$ at different $De$. At small $Oh_{p}$, the jet radius remains constant and close to the corresponding Newtonian liquid value (figure 12a: $Oh_{p}=0,0.005$). As $Oh_{p}$ increases, the jet width decreases sharply, vanishing at $Oh_{p,c}\approx 0.017$ (figure 12a: $Oh_{p}=0.013,0.018$), where $Oh_{p,c}$ marks the lower bound in $Oh_{p}$ for bubble entrainment inside the liquid (figure 12a: $Oh_{p}=0.018,0.025$ Gordillo & Rodríguez-Rodríguez (2019); Blanco-Rodríguez & Gordillo (2020); Rodríguez-Díaz et al. (2023)). Interestingly, the prediction for Newtonian liquids applies well to viscoelastic liquids by substituting $Oh_{p}$ for $Oh_{s}$ (figure 12b). Beyond $Oh_{p,c}$, the jet radius becomes ill-defined as the jet gradually widens (figure 12a: $Oh_{p}=0.035$), first reaching the dropping transition at $Oh_{p,d}\approx 0.048$ (figures 7 and 12b) and ultimately vanishing at $Oh_{p,j}\approx 0.129$ (figure 8b).
5 Conclusion and outlook
This work elucidates the effects of viscoelasticity on Worthington jet formation and droplet ejection, by contrasting it with Newtonian fluid behavior. The process is governed by two key dimensionless parameters: the elastocapillary number $Ec$, comparing elastic and capillary forces, and the Deborah number $De$, relating the relaxation time of the polymeric liquid to the inertiocapillary timescale. We identify three distinct regimes in viscoelastic media, analogous to Newtonian fluids: (i) jet formation with droplet ejection, (ii) jets without droplets, and (iii) complete jet suppression. However, the transitions between these regimes now depend on $Ec$ and $De$ rather than solely on the solvent Ohnesorge number $Oh_{s}$. Notably, while viscoelasticity significantly alters jet dynamics, it does not affect the capillary wave speed.
Analysis across the $Ec$-$De$ phase space reveals markedly different behaviors in two limiting cases. For polymeric liquids with permanent memory ($De\to\infty$), transitions occur at fixed $Ec$, independently of $De$. The jetting transition $Ec_{j}$ is independent of $Oh_{s}$, while the dropping transition $Ec_{d}$ exhibits strong $Oh_{s}$ dependence. Remarkably, this infinite $De$ asymptote extends down to $De\approx 1$ due to the finite process timescale. Conversely, for polymeric liquids with poor memory ($De\to 0$), both transitions occur at constant polymeric Ohnesorge number $Oh_{p}=Ec\times De$, indicating that addition of polymers introduces an excess viscous stress in this limit. These transitions are independent of $Oh_{s}$. Using a slender jet approach (Driessen et al., 2013; Gordillo et al., 2020; Eggers & Fontelos, 2015), we provide further insights into these transitions, examining the competition between elastic stresses and inertial flow focusing that governs jet formation and droplet ejection. This analysis helps to explain the observed scaling behaviors and transition criteria.
Our findings have important implications for understanding and controlling bubble bursting in viscoelastic fluids, with relevance to biological processes (Walls et al., 2017), such as airborne disease transmission (Bourouiba, 2021), and industrial applications, such as inkjet printing (Lohse, 2022). The results highlight how polymer additives can dramatically alter spray formation, potentially enhancing or suppressing aerosol production depending on the fluid properties (Kant et al., 2023). This work also opens several avenues for future research. Further investigation is needed into the universal behavior near the infinite $De$ asymptote, including the development of closed-form $De$ expansions for free surface flows (Sen et al., 2021; França et al., 2024; Sen et al., 2024; Balasubramanian et al., 2024; Boyko et al., 2024; Hinch et al., 2024). The mechanism underlying the $Oh_{s}$ sensitivity of transition $Ec$ values at high $De$ requires further clarification. Additionally, extending our analysis to nonlinear viscoelastic models would provide valuable insights into the role of shear-thinning behavior and finite extensibility on bursting bubbles, addressing limitations of the current model (Zinelis et al., 2023; McKinley & Sridhar, 2002; Snoeijer et al., 2020). This approach would allow quantification of discrepancies between experiments and simulations, often attributed to inherent issues with the Oldroyd-B model, thereby enhancing our understanding of viscoelastic jets (Gaillard et al., 2024). Indeed, the numerical method developed here, freely available at Sanjay & Dixit (2024), provides a generalized framework readily adaptable to any model within the Oldroyd-B family of upper convective derivative models (Snoeijer et al., 2020). Furthermore, as higher Bond numbers are observed in many scenarios (Walls et al., 2015; Ghabache et al., 2014; Deike et al., 2018; Krishnan et al., 2017), exploring their combined effect with viscoelasticity on the overall dynamics would provide valuable insights into such experiments (Rodríguez-Díaz et al., 2023). Lastly, studying interactions of multiple bubbles (Singh & Das, 2019) at the liquid-gas free surface will provide further insights into pathogen transport.
Extensions of this work could also explore coated bubbles (Dollet et al., 2019; Yang et al., 2023) or those with surface elasticity (Ji et al., 2023), and incorporate surfactants that alter bulk or interfacial properties (Constante-Amores et al., 2021; Lohse, 2022; Pierre et al., 2022; Pico et al., 2024). Utilizing the current numerical framework to investigate the effects of bubble motion (Beris et al., 1985; Moschopoulos et al., 2021) and oscillations in viscoelastic media (Dollet et al., 2019; Oratis et al., 2024) on overall dynamics before bursting would also be beneficial. This model provides a general framework for studying both Newtonian viscous and non-Newtonian elastic effects. As a future perspective, it would be worthwhile to study phenomena such as wrinkling (Debrégeas et al., 1998; Oratis et al., 2020; Davidovitch & Klein, 2024) and buckling (Le Merrer et al., 2012; Timoshenko & Gere, 2012), which occur in various viscoelastic systems (Schmalholz & Podladchikov, 1999; Lee & Dalnoki-Veress, 2024; Matoz-Fernandez et al., 2020). By encompassing both viscous and elastic behaviors, this approach enables a comprehensive study of these interconnected instabilities, elucidating their underlying mechanisms and relationships as envisioned by Stokes (1845); Lord Rayleigh (1896); Taylor (1969). Moreover, integrating viscoelastic and elastoviscoplastic (França et al., 2024; Balasubramanian et al., 2024) properties into recently developed analytical methods for capillary wave propagation and convergence, such as those by Kayal et al. (2024), could yield a deeper theoretical understanding of the phenomenon.
In conclusion, this study provides a comprehensive characterization of bubble bursting in viscoelastic media, interpreting the interplay between elastic, viscous, and capillary forces by moving in the $Oh_{s}$-$Ec$-$De$ phase space. The insights gained advance our fundamental understanding of free-surface flows in complex fluids and pave the way for more precise control of spray formation in various applications.
Supplementary data. Supplementary material and movies are available at xxxx
Code availability. The codes used in the present article are permanently available at Sanjay & Dixit (2024).
Acknowledgments. We would like to thank Jacco Snoeijer, Maziyar Jalaal, Uddalok Sen, Andrea Prosperetti, and Vincent Bertin for the discussions.
Funding. We acknowledge the funding from the MIST consortium. This publication is part of the project MIST with project number P20-35 of the research programme Perspectief, which is (partly) financed by the Dutch Research Council (NWO). We also acknowledge the NWO-Canon grant FIP-II grant. This work was carried out on the national e-infrastructure of SURFsara, a subsidiary of SURF cooperation, the collaborative ICT organization for Dutch education and research. This work was sponsored by NWO - Domain Science for the use of supercomputer facilities.
Declaration of Interests. The authors report no conflict of interest.
Appendix A The Newtonian limit of bursting bubble dynamics
Figure A 1: Temporal evolution of bubble cavity collapse in Newtonian liquid for $Oh_{s}=$ (a) $0.0025$, (b) $0.02$, and (c) $0.1$. The left panel represents the magnitude of the velocity field normalized by the inertiocapillary velocity, while the right panel shows the local viscous dissipation on a $\log_{10}$ scale. See also the supplementary movies SM3.
The dynamics of bursting bubbles in Newtonian media are solely dictated by the Ohnesorge number $Oh_{s}$ in the limit of very small bubbles (Bond number $Bo\ll 1$). Figure A 1 illustrates representative cases at varying $Oh_{s}$ for $Bo=0.001$. At low $Oh_{s}$, capillary waves propagate along the cavity, converging at its base to form a Worthington jet that subsequently fragments into droplets (figure A 1a). In this limit, multiple undamped capillary waves collide at the cavity’s bottom, generating a thick Worthington jet. Increasing $Oh_{s}$ dampens short-wavelength capillary waves, allowing the dominant wave to focus more effectively and produce a thinner jet. This explains the observed decrease in jet width with increasing $Oh_{s}$ (Gordillo & Blanco-Rodríguez, 2023), until a critical $Oh_{c}\approx 0.03$ (at $Bo=0.001$) where the jet becomes extremely narrow, approaching a singularity (Blanco-Rodríguez & Gordillo, 2020). Concurrently, the size of the first ejected droplet diminishes with increasing $Oh_{s}$ (Gordillo & Rodríguez-Rodríguez, 2019). As $Oh_{s}$ further increases, bubble entrainment occurs. Beyond $Oh_{s,d}=0.0375$, vertical droplet ejection ceases; instead, the jet undergoes Rayleigh–Plateau instability, producing droplets that fall back into the pool (Blanco-Rodríguez & Gordillo, 2020; Walls et al., 2015; Deike et al., 2018). As $Oh_{s}$ increases ($Oh_{s}>0.045$), viscous dissipation becomes more prominent, resulting in jet formation without droplet ejection (figure A 1b). Further increase in $Oh_{s}$ beyond $Oh_{s,j}=0.11$ completely suppresses jet formation (figure A 1c, also see Sanjay et al. (2022)).